OpenStudy (anonymous):

Celine has a bottle that contains 20% of milk and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of milk and water in the bottle. Define the variable used in the equation. (5 points) Part B: How many liters of milk are present in the bottle? Show your work. (5 points)

OpenStudy (anonymous):

keep in mind that Celine has a bottle, 20% of that is milk, the rest, 80%, is water the bottle has a total of 1liter of water, meaning that 80% of the bottle is 1liter because 80% of the bottle, the rest, is water

OpenStudy (anonymous):

ok... so we know that 20% is milk and that 80% is water, and that 80% is 1liter so liters1x%8020⟹1x=8020

OpenStudy (phi):

let x be the total amount of liquid (water and milk) (measured in liters) we know 80% of x is 1 liter 0.8x= 1 x = 1/0.8 = 1.25 so the total amount of liquid is 1.25 liters for part b) 20% of x is milk 0.2x = milk 0.2*1.25 = milk 0.25 liters is milk